OpenStudy (lexiquesse):

use natural log to solve 6e^(4x) -2=3

OpenStudy (lexiquesse):

@zepdrix could you help?

zepdrix (zepdrix):

So let's try to isolate the exponential e^(stuff) before applying any log business. Do you understand how to apply those steps? :) How to isolate the e^(stuff)

OpenStudy (lexiquesse):

nope

zepdrix (zepdrix):

\(\large\rm 6e^{4x}-2=3\) To undo the subtraction, we'll add 2 to each side, \(\large\rm 6e^{4x}=5\) To undo the 6 multiplying, we'll divide both sides by 6, ya?

OpenStudy (lexiquesse):

okay, next?

zepdrix (zepdrix):

So we have,\[\large\rm e^{4x}=\frac56\]

zepdrix (zepdrix):

Now that our exponential is isolated (alone), we can apply our natural log to each side, which will hopefully help us get the x out of the exponent position like it did in the previous problem.

zepdrix (zepdrix):

\[\large\rm \ln\left(e^{4x}\right)=\ln\left(\frac56\right)\]

zepdrix (zepdrix):

Remember the rule that will help us here?

OpenStudy (lexiquesse):

is it something to do with taking the ^4x out of the ln? 4xln e=ln(5/6)

zepdrix (zepdrix):

Mmmm ok good. The 4x comes down in front of the log.\[\large\rm 4x \ln e=\ln\left(\frac56\right)\]

zepdrix (zepdrix):

Recall that natural log is a log with base value e,\[\large\rm \ln(x)=\log_e(x)\] Whenever the `base of the log` matches the `contents(argument) of the log`, then it simplifies to 1. Example: \(\large\rm log_4(4)=1\) Do you see how this will help us clean up the left side of our equation?

OpenStudy (lexiquesse):

yes, so lne=1

zepdrix (zepdrix):

\[\large\rm 4x=\ln\left(\frac56\right)\]k great.

OpenStudy (lexiquesse):

divide ln(5/6) by 4?

zepdrix (zepdrix):

Mmm good good good, that will be your final step!

OpenStudy (lexiquesse):

-.046?

zepdrix (zepdrix):

Yay good job!

OpenStudy (lexiquesse):

awesome!!! thankyou sooo much! do you mind answering a few more questions? its for a final review worksheet. its okay if you dont want to

zepdrix (zepdrix):

lets try another one :) I might need a break after that

OpenStudy (lexiquesse):

these are some that its okay if you don't understand if you do could we do one? use a table to solve 6^4x =63 rewrite y= square root 9x-36 =4 to make it easy to graph using a translation

zepdrix (zepdrix):

So we're going to choose some x values, and try to end up with 63. Let's start with x=2 how bout. What do you get for 6^(4x)?

OpenStudy (lexiquesse):

1679616

zepdrix (zepdrix):

Ahhh ok so we need a much much much smaller x value then ^^ my bad. It's going to be smaller than 1 even.